OpenStudy (anonymous):

CALCULUS II: I need to find the Maclaurin series of cos^2 x by using the Maclaurin series of cos x and a half angle formula. Any idea how to do that?

OpenStudy (anonymous):

@KingGeorge

OpenStudy (kinggeorge):

A half-angle formula?

OpenStudy (anonymous):

Yeah, like cos^2 x = (1 + cos 2x)/2

OpenStudy (kinggeorge):

Right. I was just going to suggest version of that.

OpenStudy (kinggeorge):

Anyways, can you show me what the MacLaurin series for \(\cos(x)\) is?

OpenStudy (anonymous):

Sure its\[\cos x = 1 - \frac{ x^2 }{ 2!}+\frac{ x^4 }{ 4!} -\frac{ x^6 }{ 6!}+...+(-1)^n\frac{ x^(2n) }{ (2n)!}\] the x(2n) is supposed to be x^(2n). For some reason it won't go up to the exponent place.

OpenStudy (kinggeorge):

You need to use { and } instead of ( and ) to get the 2n in the exponent. But yes, that's correct. Now, if you look at your half-angle formula, you'll notice it has a \(\cos(2x)\). To get the series for that, simply plug in \(2x\) for \(x\) in the series you just wrote out.

OpenStudy (kinggeorge):

You should get \[\cos (2x) = 1 - \frac{ (2x)^2 }{ 2!}+\frac{ (2x)^4 }{ 4!} -\frac{ (2x)^6 }{ 6!}+...+(-1)^n\frac{ (2x)^{2n} }{ (2n)!}\]

OpenStudy (anonymous):

But how do you deal with the /2 part of the half angle formula?

OpenStudy (kinggeorge):

We're getting there. When dealing with series, it's best to do everything in simple steps.

OpenStudy (anonymous):

Okay, then would be next?

OpenStudy (kinggeorge):

Next, we have the \(1+\cos(2x)\) in the numerator. So we add 1 to the series.\[1+\cos (2x) = 2 - \frac{ 2^2x^2 }{ 2!}+\frac{ 2^4x^4 }{ 4!} -\frac{ 2^6x^6 }{ 6!}+...+(-1)^n\frac{2^{2n}x^{2n} }{ (2n)!}\]

OpenStudy (kinggeorge):

Finally, you divide everything by 2.

OpenStudy (kinggeorge):

And the simplest way to write the series after that, is just to take one off of the power on top of the 2's.\[\frac{1+\cos (2x)}{2} = 1 - \frac{ 2x^2 }{ 2!}+\frac{ 2^3x^4 }{ 4!} -\frac{ 2^5x^6 }{ 6!}+...+(-1)^n\frac{2^{2n-1}x^{2n} }{ (2n)!}\]

OpenStudy (anonymous):

Oh!!! okay. Thank you so much!!!

OpenStudy (kinggeorge):

You're welcome.